# MasonTigers at SemEval-2024 Task 1: An Ensemble Approach for Semantic Textual Relatedness

### v1

@@ 7 Conclusion
We experimented with diverse methodologies on the dataset provided by the organizers, including statistical machine learning approaches, transformer-based models, language-specific BERTs, and sentence BERT. In the supervised task, with no restrictions on the model or data, we utilized the available training dataset. Conversely, the unsupervised task, lacking training data, presented challenges, leading us to use language-specific BERTs and statistical machine learning approaches. The cross-lingual task imposed more stringent restrictions, requiring us to use training data from other languages in Track A, excluding the target language. In addition to statistical ML models, we integrated language-specific BERTs closely aligned with the geography and culture of the target language, as the use of LLMs was constrained due to unknown training data.
Our ensemble approach exhibited superior performance compared to individual model experiments. However, the task’s inherent difficulty became evident in instances where relatively small datasets presented challenges for effective model learning.
Limitations
Semantic textual relatedness tasks face challenges like subjectivity, context dependency, and ambiguity due to multiple meanings and cultural differences. Limited data, domain specificity, short texts, and biases hinder accuracy. Ongoing research is crucial to address these limitations and improve model accuracy.
Appendix A Result of Supervised Track
@@
Table 4: Results for Supervised Track (EN : ElasticNet, LR : Linear Regression, SC : Spearman Correlation)
Appendix B Result of Unsupervised Track
@@
Table 5: Results for Unsupervised Track (EN : ElasticNet, LR : Linear Regression, SC : Spearman Correlation)
Appendix C Result of Cross-Lingual Track
@@
Table 6: Results for Cross-Lingual Track (EN : ElasticNet, LR : Linear Regression, SC : Spearman Correlation)

### v2

@@ 7 Conclusion
We experimented with various methodologies on the dataset provided by the organizers, including statistical machine learning approaches, transformer-based models, language-specific BERTs, and sentence BERT. In the supervised task (Track A), with no restrictions on the model or data, we utilized the available training dataset. Conversely, the unsupervised task (Track B), lacking training data, presented challenges, leading us to use language-specific BERTs and statistical machine learning approaches. The cross-lingual track (Track C) imposed more stringent restrictions, requiring us to use training data from other languages in Track A, excluding the target language. In addition to statistical ML models, we integrated language-specific BERTs closely aligned with the geography and culture of the target language, as the use of LLMs was constrained due to unknown training data.
We show that our ensemble approach exhibited superior performance compared to individual model experiments. However, the task’s inherent difficulty became evident in instances where relatively small datasets presented challenges for effective model learning. Semantic textual relatedness tasks face challenges like subjectivity, context dependency, and ambiguity due to multiple meanings and cultural differences. Limited data, domain specificity, short texts, and biases hinder accuracy. Ongoing research is crucial to address these limitations and improve model accuracy.
Acknowledgements
We would like to thank the shared task organizers for providing participants with the dataset used in this paper.
Appendix A Appendix
@@
Table 4: Results of Track A (Supervised) (EN : ElasticNet, LR : Linear Regression, SC : Spearman correlation)
@@
Table 5: Results for Track B (Unsupervised) (EN : ElasticNet, LR : Linear Regression, SC : Spearman Correlation)
@@
Table 6: Results for Track C (Cross-lingual) (EN : ElasticNet, LR : Linear Regression, SC : Spearman Correlation)